outline the steps you would use to prepare a 10% salt solution from a 25% salt solution for a 100 ml.
Answer by stanbon(75887)

outline the steps you would use to prepare a 10% salt solution from a 25% salt solution for a 100 ml.
------------------
Equation:
salt + salt = salt
0.25x + 0(100-x) = 0.10(100 ml)
Multiply both sides by 100 to get:
25x = 10(100)
x = 1000/25 = 40 ml (amount of 10% solution you need in the mix)
100-x = 60 ml (amount of water you need in the mix)
=======================================================
Cheers,
Stan H.